# Burnside type results for fusion categories

### v10

@@ 12.2. Dual-Burnside property
We made the abbreviation (A)RN for (abelian) real non-negative. Note that [LPW21, Corollary 8.5] means that the Grothendieck ring, if commutative, of a unitary fusion category, has a RN dual (hypergroup).
Question 12.1.
@@
Recall that a commutative fusion ring $F$ with RN dual is called dual-Burnside when a column of its character table has a zero entry if and only if the squared norm of the column (i.e. the corresponding formal codegree) has value strictly less than $\mathrm{FPdim}(F)$, or alternatively, when $P^{2}$ is an idempotent, where
$P=\prod_{i\in I}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
and $\{x_{i}\}_{i\in I}$ is the set of basis elements (Corollary 4.16). We warn the reader that for a non-abelian group $G$, the product of all its elements depends on the ordering, in fact, the set of all possible products forms a $G^{\prime}$-coset, see [DH82].
Recall that [INW99, Theorem B] states that for all finite nilpotent group $G$ then the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside. It is extended to every nilpotent dualizable ARN-hypergroup by Theorem 1.12. So every commutative nilpotent fusion ring with RN dual is both Burnside and dual-Burnside. Note that all the commutative nilpotent fusion rings we checked have RN dual.
Question 12.2.
Is there a commutative nilpotent fusion ring without RN dual?
Let us call a finite group $G$ dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside. So a nilpotent finite group is dual-Burnside. There are exactly $144$ finite groups of order less than $32$, and among them, exactly $30$ ones are non-nilpotent, and among them, only $\mathrm{SL}(2,3)$ is dual-Burnside (so the $29$ other ones, e.g. $S_{3}$ and $D_{5}$, are not dual-Burnside).
Observe that a finite group $G$ is centerless if and only if $K(\mathrm{Rep}(G))$ has a perfect dual (see Definition 2.39, Theorem 7.8 and Lemma 11.4). So a centerless finite group is dual-Burnside if and only if every non-$\mathrm{FPdim}$ column of its character table has a zero entry. We checked that every centerless and dual-Burnside finite group $G$ of order $|G|\leq 1000$ is almost simple (i.e. $S\subseteq G\subseteq\mathrm{Aut}(S)$ with $S$ non-abelian simple); but $G=A_{5}\times A_{5}$, of order $3600$, is not almost simple. A finite non-abelian simple group is centerless, so by [Pal] and its answers, all the non-alternating finite simple groups are dual-Burnside, except the Mathieu groups $M_{22}$ and $M_{24}$. The alternating group $A_{n}$, $5\leq n\leq 19$, is dual-Burnside if and only if $n\not\in\{7,11,13,15,16,18,19\}$. Recall that a finite group $G$ is simple if and only if $K(\mathrm{Rep}(G))$ is simple.
@@ 12.3. Modular categories
Let us provide some applications of Theorem 1.16. There are exactly $43$ different types for a $\frac{1}{2}$-Frobenius integral fusion ring up to rank $11$ (see [ABPP]), and among them, exactly $7$ ones (below) are excluded from modular categorification by Theorem 1.18.
$[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],$
$[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4].$
@@
They are excluded by Theorem 1.18 from modular categorification, because $3$ is a powerless factor of the $\mathrm{FPdim}$, whereas $3$ does not divide $4$, the number of grouplike elements. Note that the above fusion data $3$ and $4$ can be generalized as follows. Let $G$ be a group of order $n^{2}$, and let $K$ be a group of order $m+1$. Consider the following $\frac{1}{2}$-Frobenius integral fusion ring:
@@
It is excluded by Theorem 1.16 (1.13) from modular categorification if $n>1$ and $m+1$ has a prime factor not dividing $n$.
Availability of data and materials
The datasets generated during and/or analyzed during the current study are available from the second author on reasonable request.
Declaration
Conflict of interests: The authors declare that they have no conflict of interest.

### v11

@@ 12.2. Dual-Burnside property
We abbreviate “(A)RN” for “(abelian) real non-negative.” According to [LPW21, Corollary 8.5], the Grothendieck ring of a unitary fusion category, if commutative, has an RN dual (hypergroup).
Question 12.2.
@@
Recall that a commutative fusion ring $F$ is called dual-Burnside if a column of its character table has a zero entry if and only if the squared norm of the column (the corresponding formal codegree) is strictly less than $\mathrm{FPdim}(F)$. When $F$ has an RN dual, it is equivalent to $P^{2}$ being idempotent, where
$P=\prod_{i\in\mathcal{I}}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
and $\{x_{i}\}_{i\in I}$ is the set of basis elements (see Corollary 4.16). Note that for a non-abelian group $G$, the product of all its elements depends on the ordering; in fact, the set of all possible products forms a $G^{\prime}$-coset (see [DH82]).
According to [INW99, Theorem B], for any finite nilpotent group $G$, the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside. This extends to every nilpotent dualizable ARN-hypergroup by Theorem 1.15. Thus, every commutative nilpotent fusion ring with an RN dual is both Burnside and dual-Burnside. All commutative nilpotent fusion rings we examined have RN duals.
Question 12.3.
Is there a commutative nilpotent fusion ring without an RN dual?
We define a finite group $G$ as dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside. Therefore, a nilpotent finite group is dual-Burnside. There are exactly 144 finite groups of order less than 32, among which 30 are non-nilpotent. Of these, only $\mathrm{SL}(2,3)$ is dual-Burnside, while the other 29, such as $S_{3}$ and $D_{5}$, are not.
[…]
Remark 12.4.
By Theorem 1.8, a finite group $G$ is dual-Burnside if and only if
$\left(\prod_{\chi\in\mathrm{Irr}(G)}\frac{\chi}{\chi(1)}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{\chi\in\mathrm{Irr}(G/Z(G))}\chi(1)\chi\right).$
@@ 12.3. Modular categories
Let us present some applications of Theorem 1.19. There are exactly 71 distinct types of half-Frobenius integral fusion rings with ranks up to 12 (see [ABPP]). Of these, 15 types (listed below) are excluded from modular categorification by Theorems 1.19 or 1.22.
[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[ 1 , 1 , 1 , 1 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 ] ,
[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] ,
[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 , 4 ] ,
[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 8 , 8 ] , [ 1 , 1 , 2 , 2 , 2 , 2 , 3 , 3 , 6 , 6 , 6 , 6 ] .
@@
They are excluded from modular categorification by Theorem 1.19 (1.14) because $\mathcal{V}(\mathcal{C})=\mathcal{V}(12)=\{2,3\}$, while $3\not\in\mathcal{V}(d_{i})$ for any $i$, and $\mathcal{V}(\mathcal{C}_{pt})=\{2\}$. Additionally, Theorem 1.22 rules them out since 3 is a (powerless) factor of $\mathrm{FPdim}(\mathcal{C})$, but does not divide $\mathrm{FPdim}(\mathcal{C}_{pt})=4$.
Note that the fusion data for cases 3 and 4 can be generalized. Consider a group $G$ with order $n^{2}$ and a group $K$ with order $m+1$. Then, analyze the following half-Frobenius integral fusion ring:
@@
According to Theorem 1.19 (1.14), it is excluded from modular categorification if $n>1$ and $m+1$ has a prime factor that does not divide $n$.
Note that the types [1,1,1,1,1,1,1,1,1,3,3,3][1,1,1,1,1,1,1,1,1,3,3,3][ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] and [1,1,1,1,2,2,2,2,2,2,2,2][1,1,1,1,2,2,2,2,2,2,2,2][ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] are excluded from modular categorification by Theorem 1.19, but not by Theorem 1.22, since their $\mathrm{FPdim}=2^{2}3^{2}$ have no powerless prime factor.
Availability of data and materials. Data for the computations in this paper are available on reasonable request from the authors.
Conflict of interest statement. On behalf of all authors, the corresponding author declares that there are no conflicts of interest.